# FS-DFM: Fast and Accurate Long Text Generation with Few-Step Diffusion Language Model

### v3

@@ Abstract
Autoregressive language models (ARMs) generate sequences by predicting the next token conditioned on the observed prefix, and have achieved remarkable success (Yang et al., 2025; Team et al., 2023; Grattafiori et al., 2024; Li et al., 2025a). In contrast, Diffusion Language Models (DLMs) synthesize text through iterative refinement, providing stable likelihood-grounded training objectives, inherent parallelism across positions, and enhanced controllability through access to global context, capabilities that help mitigate exposure bias and naturally accommodate structure beyond left-to-right generation (Nie et al., 2025b; Austin et al., 2021). Despite these complementary advantages, both paradigms face fundamental bottlenecks. ARMs are constrained by sequential decoding, requiring one forward pass per token, which limits throughput and enforces a unidirectional dependency that hinders tasks such as reversal or order-invariant reasoning (Zhao et al., 2025; Schulman et al., 2022; Gong et al., 2025b). Although DLMs parallelize over positions, they typically require tens to hundreds of refinement steps to achieve competitive quality, effectively replacing the autoregressive depth of a single long pass with a stack of iterative model evaluations (Chen & Zhou, 2023; Arriola et al., 2025; Nie et al., 2025a). For example, to march the ARM generation quality, LLaDA requires roughly one inference step per output token (Nie et al., 2025b).
@@
Continuous-space flow-matching models have emerged as a streamlined alternative to conventional score-based diffusion trained around reverse Stochastic Differential Equations (SDEs). Instead of modeling a stochastic reverse process, flow-matching directly trains a continuous normalizing flow via vector-field regression, enabling deterministic probability-flow Ordinary Differential Equation (ODE) sampling. This approach often achieves comparable or superior likelihoods and sample quality with fewer function evaluations and more stable optimization (e.g., through optimal transport or straight-line paths) than SDE-based pipelines (Lipman et al., 2023; Song et al., 2021; Li et al., 2025b; Gong et al., 2023). Discrete Flow-Matching (DFM) extends these advantages to discrete domains by adapting flow-matching to tokenized data, providing flexible probability paths and strong perplexity at scale. However, DFM still requires a substantial number of evaluations along each trajectory, hindering applicability (Gat et al., 2024).
@@
FS-DFM builds on two pillars. First, we enrich the model with the shortcut principle (Frans et al., 2025) together with a self-consistency constraint: the network is conditioned not only on masked tokens but also on the desired step size, and trained so that a single large step agrees with the composition of multiple smaller steps. We instantiate the shortcut teacher using a Runge–Kutta formulation —we evaluate RK-2 (Heun) and classical RK-4 as ODE estimators— and adopt the one that offers the best accuracy–stability trade-off for large single/few-step updates. Second, a DFM reformulation that models sequences as a Continuous-Time Discrete Markov Chain (CTMC, Campbell et al. (2022)) and a probability velocity is learned to transport a source distribution to the data distribution along a probability path (Gat et al., 2024; Campbell et al., 2024). In this framework, a velocity field advances samples “locally” from $p_{t}$ to $p_{t+h}$. Instead, FS-DFM learns a Cumulative Scalar parameterized by the step size $h$ to enable large and reliable jumps along the probability path. The step-budgeted perspective retains the core benefits of flow-matching: stable training, parallel refinement, and bidirectional context, while elevating few-step generation to a first-class design objective.
@@ 2 Related Work
Diffusion models have demonstrated remarkable effectiveness in continuous domains such as images, achieving state-of-the-art quality across diverse tasks (Karimi Monsefi et al., 2025; Zhang et al., 2023; Navard et al., 2024; Rombach et al., 2022). Extending diffusion-style generative modeling to text, Diffusion Language Models (DLMs), have shown impressive task performance coupled with fast inference through parallel generation (Nie et al., 2025b; Ye et al., 2025; Labs et al., 2025). DLMs can be broadly grouped into continuous models, which denoise in embedding or latent space, and discrete models, which operate directly over tokens. Both categories exhibit the key advantages of diffusion, including bidirectional context, controllability, and parallel refinement (Li et al., 2025b).
@@
Continuous DLMs (Li et al., 2022; Lin et al., 2023; Lovelace et al., 2023) cast text generation as iterative denoising in a continuous embedding or latent space. A forward corruption process perturbs token embeddings or hidden states, while a parameterized denoiser reconstructs clean representations conditioned on the context. Continuous DLMs take advantage of benefits that arise naturally in continuous representations: flexible conditioning and guidance, bidirectional contextualization during refinement, and compatibility with large-scale pretraining or task-specific fine-tuning. Moreover, continuous DLMs benefits from research in diffusion models for other modalities such as images or audio as methods easily port between domains.
@@
Discrete DLMs, on the other hand, model token dynamics either through discrete-time noise schedules or continuous-time Markov formulations. DLMs such as LLaDA (Nie et al., 2025b) and Dream (Ye et al., 2025) match autoregressive models in task performance if allowed sufficient refinement steps. Effectively, they tie step count to sequence length and generation quality (Austin et al., 2021; Gong et al., 2025a; Zheng et al., 2024; Shi et al., 2024; Ou et al., 2025). One line of research accepts this high step count requirement and focuses on speeding up inference through approaches such as confidence-aware parallel decoding, tweaks to re-enable the KV Cache, and restricted attention (Wu et al., 2025; Arriola et al., 2025; Chen et al., 2025b). Another line of research focuses on reducing the number of inference steps through one-/few-step approaches (Gat et al., 2024). Our work falls within this second line of research.
@@
Prior work on one-step language generation primarily focused on continuous-space diffusion models, operating over embeddings rather than discrete tokens. DLM-One (Chen et al., 2025a) demonstrated large speedups on short sequence tasks but did not address long sequence modeling. Similarly, FlowSeq (Hu et al., 2024) operates on continuous embeddings, bypassing token-level likelihoods. Within the discrete family, recent methods typically attain quality with thousands of refinement steps, often scaling with sequence length (Gat et al., 2024; Zhao et al., 2025). SDTT (Deschenaux & Gulcehre, 2025) distills discrete DLMs for faster sampling, yet still uses 16–256 steps and focuses on short-form tasks. In contrast, our method (FS-DFM) operates in the discrete space and explicitly targets the few-step regime (1-8 steps). FS-DFM can match 1 024-step discrete-flow baselines in just 8 steps while preserving text quality.
@@ 3 Preliminaries and Background
We provide a brief summary on flow-matching, following Gat et al. (2024). For the notation used and for details on the theory behind discrete flow-matching, readers are referred to Appendix˜A.
@@
Of course, this depends on the choice of probability path $(p_{t})_{0\leq t\leq 1}$. Following Gat et al. (2024), start with
@@
To train this model, we need an appropriate loss function. We use the Bregman divergence for the loss function (Lipman et al., 2024, Equation 7.31). Starting from the velocity formulation in Equation˜3.5, we can construct a loss that encourages the model to correctly predict $p_{1|t}$. Given a sample trajectory where $x_{0}\sim p_{0}$, $x_{1}\sim p_{1}$ are related through $\pi$, and $x_{t}$ is sampled from $p_{t}(x|x_{0},x_{1})$, the per-token loss at position $i$ is:
@@ 4 Method
Our approach comprises two components: Step-Aware Discrete Flow-Matching and a Cumulative Scalar update. The first component exposes the step budget $h$ as an explicit control signal within the DFM generator $u_{t}$ and distills from a shortcut teacher model (Frans et al., 2025), allowing the learner to make large, single/few-step moves that approximate the cumulative effect of many small updates. The second is using a cumulative scalar to aid the model in jumping from a source token to a target token with large single/few-step predictions.
@@ 5.1 Experimental Setup
Step-aware DFM training incurs additional model evaluations per batch (for shortcut RK-2/RK-4 teachers), making from-scratch optimization expensive. We therefore adopt a pretrain $\rightarrow$ fine-tune protocol: first pretrain a plain DFM backbone, then fine-tune it with step-aware objectives and the cumulative scalar update. We pretrain the DFM model following Gat et al. (2024) and cover two source distributions—uniform and mask—and three model sizes: 0.169B, 1.3B, and 1.7B parameters for the uniform source, plus a 0.169B mask-source model. These checkpoints serve as initialization for FS-DFM fine-tuning. Section˜C.1 contains all architecture and training details.
We train on FineWeb-Edu (Lozhkov et al., 2024) and evaluate on WikiText-103 (Merity et al., 2017). We use GPT-2 tokenizer and, during preprocessing, we append an EOS token to each document and pack the resulting token stream into contiguous blocks of length 1 024. We concatenate shorter samples to reach the desired 1 024 token target.
@@
Measurement. We report several complementary metrics: Perplexity (PPL) measured by a fixed reference LM (gpt2-large); lower is better; and Entropy – the average uncertainty of our model’s token distributions; lower indicates sharper, more decisive predictions; Token accuracy: the fraction of model prediction that match the ground-truth; higher is better; and MAUVE (Pillutla et al., 2021), divergence-based metric that measures how similar the distribution of text generated by a model is to that of real human text. All metrics are computed on the evaluation split and averaged across sequences.
Baselines. We compare FS-DFM to a broad set of discrete diffusion and few-step generative models. In the main text, we report results against two diffusion language models, LLaDA-8B (Nie et al., 2025b) and Dream-7B (Ye et al., 2025), each in Base and Instruct variants, as well as discrete flow-matching (DFM) baselines at three model sizes (0.169B, 1.3B, 1.7B). To further contextualize performance, Section˜E.3 expands this comparison to include multi-round refinement models such as SDTT (Deschenaux & Gulcehre, 2025), hybrid discrete–continuous diffusion models such as HDLM (Fathi et al., 2025), and several masked-diffusion / re-masking systems (MDLM (Sahoo et al., 2024), SEDD (Lou et al., 2023), ReMDM (Wang et al., 2025)), all evaluated under a fixed few-step generation budget.
@@ 5.2 Results
Table 2: FS-DFM vs. diffusion LMs across step budgets. Each method receives a 512-token prefix; metrics are computed on the 512-token continuation only (ppl = perplexity, ent = entropy, MVE = MAUVE Pillutla et al. (2021)). FS-DFM attains competitive quality in few steps with stable entropy across sizes, whereas baselines generally require many steps.
@@ C.1 Pre-train model architecture
Backbone. We use a DiT-style G (Peebles & Xie, 2023) with rotary attention. Tokens are embedded once, rotary phases are computed once per forward pass and applied to $Q,K$, and each block applies multi-head self-attention followed by an MLP. A conditioning vector $c$ drives adaptive layer normalization: each block predicts per-channel shift/scale (and a small residual weight) from $c$ and applies them before attention and before the MLP. The final linear head is zero-initialized for stable start-up and produces logits; conversion of logits to a CTMC generator and ODE stepping are handled by the solver outside the backbone.
@@ E.1 Ablation Studies
Effect of source Distribution for Few-Step Generation. We study the effect of the source distribution used by DFM on few-step sampling. Following Gat et al. (2024), we compare a mask source (all positions start as a dedicated [MASK] state) against a uniform source (initialized with random tokens). With only one or a few steps, the CTMC has limited opportunity to escape [MASK]; transitions concentrate on a small set of outcomes or remain at [MASK], producing collapse and high perplexity. In contrast, the uniform source spreads initial mass over real tokens, so the model performs corrective refinements rather than first-escape jumps. This yields markedly lower perplexity with healthy entropy at tiny budgets and remains competitive as the budget grows. Table˜4 shows that uniform source is consistently superior in the few-step regime, while the performance gap narrows with larger NFE as more transitions become available. For this experiment, we used the RK-2 method for both mask and uniform sources.
@@ E.3 Expanded Baseline Comparison and Positioning of FS-DFM
SDTT (1 round) (Deschenaux & Gulcehre, 2025) 0.86 612.38 5.80
SDTT (3 rounds) (Deschenaux & Gulcehre, 2025) 0.86 376.19 5.71
SDTT (5 rounds) (Deschenaux & Gulcehre, 2025) 0.86 220.81 5.55
SDTT (7 rounds) (Deschenaux & Gulcehre, 2025) 0.86 131.71 5.26
@@
HDLM ($\epsilon=0.01$) (Fathi et al., 2025) 0.16 524 279.60 5.32
HDLM ($\gamma=0.05$) (Fathi et al., 2025) 0.16 524 192.74 5.62
@@
MDLM–OWT (Sahoo et al., 2024) 0.17 262 837.58 5.90
SEDD (Lou et al., 2023) 0.32 602.09 5.85
ReMDM–cap (Wang et al., 2025) 0.17 262 559.74 5.79
ReMDM–rescale (Wang et al., 2025) 0.17 262 560.06 5.79
ReMDM–conf (Wang et al., 2025) 0.17 262 549.86 5.79
ReMDM–loop (Wang et al., 2025) 0.17 262 1358.84 5.85
@@
A first category consists of multi-round refinement models such as SDTT (Deschenaux & Gulcehre, 2025), which improve quality through multiple decoding rounds. Although additional rounds consistently reduce perplexity, even seven rounds of refinement remain substantially behind the performance of FS-DFM at only eight sampling steps (see Table˜5). Beyond this quantitative gap, the two approaches also differ conceptually: SDTT accelerates generation by sampling from a many-step teacher and distilling its predictions, whereas FS-DFM directly learns step-size–aware finite-step Kolmogorov dynamics via RK-based ODE integration and a cumulative scalar inside the DFM framework. In other words, FS-DFM models how probability mass should evolve under different step budgets, rather than only compressing a fixed long-run sampler, which helps explain its stronger few-step behavior.
A second category includes hybrid diffusion–autoregressive models such as HDLM (Fathi et al., 2025). These models leverage continuous-time diffusion in combination with powerful sequence encoders and benefit from extremely large training sets—often exceeding 500B tokens. Despite this enormous training budget, their perplexity remains noticeably higher than that achieved by FS-DFM using a considerably smaller pre-training and fine-tuning corpus. This reinforces the training-token efficiency of our fast flow–based method.
We also include several discrete masked-diffusion and re-masking systems—ReMDM variants, and SEDD (Lou et al., 2023). These approaches attempt to improve sample quality by strategically masking tokens and applying multi-pass denoising schedules. Some of these baselines rely on exceptionally large training corpora, reaching into the hundreds of billions of tokens. Yet even under such favorable conditions, their perplexity and entropy remain less competitive. By contrast, FS-DFM achieves substantially stronger results with a dramatically smaller token budget, highlighting the impact of our distillation strategy and few-step training design.
Finally, we include the original DFM models from (Gat et al., 2024). These models are trained on the same 65.5B-token corpus as the backbone of our system, making them a direct and informative reference point. The results show that FS-DFM improves perplexity by more than a factor of four to five compared to these predecessors while preserving the same generation budget of eight sampling steps. This comparison isolates the contribution of our flow-distilled fast sampler from confounding factors such as model size or data scale.

### v4

@@ Abstract
Autoregressive language models (ARMs) generate sequences by predicting the next token conditioned on the observed prefix, and have achieved remarkable success (yang2025qwen3; team2023gemini; dubey2024llama; li2025apple). In contrast, Diffusion Language Models (DLMs) synthesize text through iterative refinement, providing stable likelihood-grounded training objectives, inherent parallelism across positions, and enhanced controllability through access to global context, capabilities that help mitigate exposure bias and naturally accommodate structure beyond left-to-right generation (nieLargeLanguageDiffusion2025; austin2021structured). Despite these complementary advantages, both paradigms face fundamental bottlenecks. ARMs are constrained by sequential decoding, requiring one forward pass per token, which limits throughput and enforces a unidirectional dependency that hinders tasks such as reversal or order-invariant reasoning (zhao2023survey; schulman2022chatgpt; gong2025diffucoder). Although DLMs parallelize over positions, they typically require tens to hundreds of refinement steps to achieve competitive quality, effectively replacing the autoregressive depth of a single long pass with a stack of iterative model evaluations (chen2023learning; arriolaBlockDiffusionInterpolating2025; nie2024scaling). For example, to march the ARM generation quality, LLaDA requires roughly one inference step per output token (nieLargeLanguageDiffusion2025).
@@
Continuous-space flow-matching models have emerged as a streamlined alternative to conventional score-based diffusion trained around reverse Stochastic Differential Equations (SDEs). Instead of modeling a stochastic reverse process, flow-matching directly trains a continuous normalizing flow via vector-field regression, enabling deterministic probability-flow Ordinary Differential Equation (ODE) sampling. This approach often achieves comparable or superior likelihoods and sample quality with fewer function evaluations and more stable optimization (e.g., through optimal transport or straight-line paths) than SDE-based pipelines (lipman2022flow; song2020score; liSurveyDiffusionLanguage2025; gong2022diffuseq). Discrete Flow-Matching (DFM) extends these advantages to discrete domains by adapting flow-matching to tokenized data, providing flexible probability paths and strong perplexity at scale. However, DFM still requires a substantial number of evaluations along each trajectory, hindering applicability (NEURIPS2024_f0d629a7).
@@
FS-DFM builds on two pillars. First, we enrich the model with the shortcut principle (frans2024one) together with a self-consistency constraint: the network is conditioned not only on masked tokens but also on the desired step size, and trained so that a single large step agrees with the composition of multiple smaller steps. We instantiate the shortcut teacher using a Runge–Kutta formulation —we evaluate RK-2 (Heun) and classical RK-4 as ODE estimators— and adopt the one that offers the best accuracy–stability trade-off for large single/few-step updates. Second, a DFM reformulation that models sequences as a Continuous-Time Discrete Markov Chain (CTMC, campbell2022continuous) and a probability velocity is learned to transport a source distribution to the data distribution along a probability path (NEURIPS2024_f0d629a7; campbell2024generative). In this framework, a velocity field advances samples “locally” from $p_{t}$ to $p_{t+h}$. Instead, FS-DFM learns a Cumulative Scalar parameterized by the step size $h$ to enable large and reliable jumps along the probability path. The step-budgeted perspective retains the core benefits of flow-matching: stable training, parallel refinement, and bidirectional context, while elevating few-step generation to a first-class design objective.
@@ 2 Related Work
Diffusion models have demonstrated remarkable effectiveness in continuous domains such as images, achieving state-of-the-art quality across diverse tasks (karimi2025taxadiffusion; zhang2023adding; navard2024knobgen; rombach2022high). Extending diffusion-style generative modeling to text, Diffusion Language Models (DLMs), have shown impressive task performance coupled with fast inference through parallel generation (nieLargeLanguageDiffusion2025; yeDream7BDiffusion2025; labsMercuryUltraFastLanguage2025). DLMs can be broadly grouped into continuous models, which denoise in embedding or latent space, and discrete models, which operate directly over tokens. Both categories exhibit the key advantages of diffusion, including bidirectional context, controllability, and parallel refinement (liSurveyDiffusionLanguage2025).
@@
Continuous DLMs (li2022diffusion; lin2023text; lovelace2023latent) cast text generation as iterative denoising in a continuous embedding or latent space. A forward corruption process perturbs token embeddings or hidden states, while a parameterized denoiser reconstructs clean representations conditioned on the context. Continuous DLMs take advantage of benefits that arise naturally in continuous representations: flexible conditioning and guidance, bidirectional contextualization during refinement, and compatibility with large-scale pretraining or task-specific fine-tuning. Moreover, continuous DLMs benefits from research in diffusion models for other modalities such as images or audio as methods easily port between domains.
@@
Discrete DLMs, on the other hand, model token dynamics either through discrete-time noise schedules or continuous-time Markov formulations. DLMs such as LLaDA (nieLargeLanguageDiffusion2025) and Dream (yeDream7BDiffusion2025) match autoregressive models in task performance if allowed sufficient refinement steps. Effectively, they tie step count to sequence length and generation quality (austin2021structured; gong2024scaling; zheng2023reparameterized; shi2024simplified; ou2024your). One line of research accepts this high step count requirement and focuses on speeding up inference through approaches such as confidence-aware parallel decoding, tweaks to re-enable the KV Cache, and restricted attention (wuFastdLLMTrainingfreeAcceleration2025; arriolaBlockDiffusionInterpolating2025; chenDPadEfficientDiffusion2025). Another line of research focuses on reducing the number of inference steps through one-/few-step approaches (NEURIPS2024_f0d629a7). Our work falls within this second line of research.
@@
Prior work on one-step language generation primarily focused on continuous-space diffusion models, operating over embeddings rather than discrete tokens. DLM-One (chen2025dlm) demonstrated large speedups on short sequence tasks but did not address long sequence modeling. Similarly, FlowSeq (hu2024flow) operates on continuous embeddings, bypassing token-level likelihoods. Within the discrete family, recent methods typically attain quality with thousands of refinement steps, often scaling with sequence length (NEURIPS2024_f0d629a7; zhao2023survey). SDTT (deschenaux2024beyond) distills discrete DLMs for faster sampling, yet still uses 16–256 steps and focuses on short-form tasks. In contrast, our method (FS-DFM) operates in the discrete space and explicitly targets the few-step regime (1-8 steps). FS-DFM can match 1 024-step discrete-flow baselines in just 8 steps while preserving text quality.
@@ 3 Preliminaries and Background
We provide a brief summary on flow-matching, following NEURIPS2024_f0d629a7. For the notation used and for details on the theory behind discrete flow-matching, readers are referred to Appendix˜A.
@@
Of course, this depends on the choice of probability path $(p_{t})_{0\leq t\leq 1}$. Following NEURIPS2024_f0d629a7, start with
@@
To train this model, we need an appropriate loss function. We use the Bregman divergence for the loss function (lipman2024flow, Equation 7.31). Starting from the velocity formulation in Equation˜3.5, we can construct a loss that encourages the model to correctly predict $p_{1|t}$. Given a sample trajectory where $x_{0}\sim p_{0}$, $x_{1}\sim p_{1}$ are related through $\pi$, and $x_{t}$ is sampled from $p_{t}(x|x_{0},x_{1})$, the per-token loss at position $i$ is:
@@ 4 Method
Our approach comprises two components: Step-Aware Discrete Flow-Matching and a Cumulative Scalar update. The first component exposes the step budget $h$ as an explicit control signal within the DFM generator $u_{t}$ and distills from a shortcut teacher model (frans2024one), allowing the learner to make large, single/few-step moves that approximate the cumulative effect of many small updates. The second is using a cumulative scalar to aid the model in jumping from a source token to a target token with large single/few-step predictions.
@@ 5.1 Experimental Setup
Step-aware DFM training incurs additional model evaluations per batch (for shortcut RK-2/RK-4 teachers), making from-scratch optimization expensive. We therefore adopt a pretrain $\rightarrow$ fine-tune protocol: first pretrain a plain DFM backbone, then fine-tune it with step-aware objectives and the cumulative scalar update. We pretrain the DFM model following NEURIPS2024_f0d629a7 and cover two source distributions—uniform and mask—and three model sizes: 0.169B, 1.3B, and 1.7B parameters for the uniform source, plus a 0.169B mask-source model. These checkpoints serve as initialization for FS-DFM fine-tuning. Section˜C.1 contains all architecture and training details.
We train on FineWeb-Edu (lozhkov2024fineweb-edu) and evaluate on WikiText-103 (merity2016pointer). We use GPT-2 tokenizer and, during preprocessing, we append an EOS token to each document and pack the resulting token stream into contiguous blocks of length 1 024. We concatenate shorter samples to reach the desired 1 024 token target.
@@
Measurement. We report several complementary metrics: Perplexity (PPL) measured by a fixed reference LM (gpt2-large); lower is better; and Entropy – the average uncertainty of our model’s token distributions; lower indicates sharper, more decisive predictions; Token accuracy: the fraction of model prediction that match the ground-truth; higher is better; and MAUVE (pillutlaMAUVEMeasuringGap2021), divergence-based metric that measures how similar the distribution of text generated by a model is to that of real human text. All metrics are computed on the evaluation split and averaged across sequences.
Baselines. We compare FS-DFM to a broad set of discrete diffusion and few-step generative models. In the main text, we report results against two diffusion language models, LLaDA-8B (nieLargeLanguageDiffusion2025) and Dream-7B (yeDream7BDiffusion2025), each in Base and Instruct variants, as well as discrete flow-matching (DFM) baselines at three model sizes (0.169B, 1.3B, 1.7B). To further contextualize performance, Section˜E.3 expands this comparison to include multi-round refinement models such as SDTT (deschenaux2024beyond), hybrid discrete–continuous diffusion models such as HDLM (fathi2025unifying), and several masked-diffusion / re-masking systems (MDLM (sahoo2024simple), SEDD (lou2023discrete), ReMDM (wang2025remasking)), all evaluated under a fixed few-step generation budget.
@@ 5.2 Results
Table 2: FS-DFM vs. diffusion LMs across step budgets. Each method receives a 512-token prefix; metrics are computed on the 512-token continuation only (ppl = perplexity, ent = entropy, MVE = MAUVE pillutlaMAUVEMeasuringGap2021). FS-DFM attains competitive quality in few steps with stable entropy across sizes, whereas baselines generally require many steps.
@@ 6 Discussion and Conclusion
Acknowledgement
We would like to thank Ruixiang Zhang for his valuable insights and continued support throughout the project, particularly for his thoughtful feedback and guidance in addressing key challenges.
@@ C.1 Pre-train model architecture
Backbone. We use a DiT-style G (Peebles2022DiT) with rotary attention. Tokens are embedded once, rotary phases are computed once per forward pass and applied to $Q,K$, and each block applies multi-head self-attention followed by an MLP. A conditioning vector $c$ drives adaptive layer normalization: each block predicts per-channel shift/scale (and a small residual weight) from $c$ and applies them before attention and before the MLP. The final linear head is zero-initialized for stable start-up and produces logits; conversion of logits to a CTMC generator and ODE stepping are handled by the solver outside the backbone.
@@ E.1 Ablation Studies
Effect of source Distribution for Few-Step Generation. We study the effect of the source distribution used by DFM on few-step sampling. Following NEURIPS2024_f0d629a7, we compare a mask source (all positions start as a dedicated [MASK] state) against a uniform source (initialized with random tokens). With only one or a few steps, the CTMC has limited opportunity to escape [MASK]; transitions concentrate on a small set of outcomes or remain at [MASK], producing collapse and high perplexity. In contrast, the uniform source spreads initial mass over real tokens, so the model performs corrective refinements rather than first-escape jumps. This yields markedly lower perplexity with healthy entropy at tiny budgets and remains competitive as the budget grows. Table˜4 shows that uniform source is consistently superior in the few-step regime, while the performance gap narrows with larger NFE as more transitions become available. For this experiment, we used the RK-2 method for both mask and uniform sources.
@@ E.3 Expanded Baseline Comparison and Positioning of FS-DFM
SDTT (1 round) (deschenaux2024beyond) 0.86 612.38 5.80
SDTT (3 rounds) (deschenaux2024beyond) 0.86 376.19 5.71
SDTT (5 rounds) (deschenaux2024beyond) 0.86 220.81 5.55
SDTT (7 rounds) (deschenaux2024beyond) 0.86 131.71 5.26
@@
HDLM ($\epsilon=0.01$) (fathi2025unifying) 0.16 524 279.60 5.32
HDLM ($\gamma=0.05$) (fathi2025unifying) 0.16 524 192.74 5.62
@@
MDLM–OWT (sahoo2024simple) 0.17 262 837.58 5.90
SEDD (lou2023discrete) 0.32 602.09 5.85
ReMDM–cap (wang2025remasking) 0.17 262 559.74 5.79
ReMDM–rescale (wang2025remasking) 0.17 262 560.06 5.79
ReMDM–conf (wang2025remasking) 0.17 262 549.86 5.79
ReMDM–loop (wang2025remasking) 0.17 262 1358.84 5.85
@@
A first category consists of multi-round refinement models such as SDTT (deschenaux2024beyond), which improve quality through multiple decoding rounds. Although additional rounds consistently reduce perplexity, even seven rounds of refinement remain substantially behind the performance of FS-DFM at only eight sampling steps (see Table˜5). Beyond this quantitative gap, the two approaches also differ conceptually: SDTT accelerates generation by sampling from a many-step teacher and distilling its predictions, whereas FS-DFM directly learns step-size–aware finite-step Kolmogorov dynamics via RK-based ODE integration and a cumulative scalar inside the DFM framework. In other words, FS-DFM models how probability mass should evolve under different step budgets, rather than only compressing a fixed long-run sampler, which helps explain its stronger few-step behavior.
A second category includes hybrid diffusion–autoregressive models such as HDLM (fathi2025unifying). These models leverage continuous-time diffusion in combination with powerful sequence encoders and benefit from extremely large training sets—often exceeding 500B tokens. Despite this enormous training budget, their perplexity remains noticeably higher than that achieved by FS-DFM using a considerably smaller pre-training and fine-tuning corpus. This reinforces the training-token efficiency of our fast flow–based method.
We also include several discrete masked-diffusion and re-masking systems—ReMDM variants, and SEDD (lou2023discrete). These approaches attempt to improve sample quality by strategically masking tokens and applying multi-pass denoising schedules. Some of these baselines rely on exceptionally large training corpora, reaching into the hundreds of billions of tokens. Yet even under such favorable conditions, their perplexity and entropy remain less competitive. By contrast, FS-DFM achieves substantially stronger results with a dramatically smaller token budget, highlighting the impact of our distillation strategy and few-step training design.
Finally, we include the original DFM models from (NEURIPS2024_f0d629a7). These models are trained on the same 65.5B-token corpus as the backbone of our system, making them a direct and informative reference point. The results show that FS-DFM improves perplexity by more than a factor of four to five compared to these predecessors while preserving the same generation budget of eight sampling steps. This comparison isolates the contribution of our flow-distilled fast sampler from confounding factors such as model size or data scale.